OpenStudy (anonymous):

solve by graphing x-y=3 2x+y=3

OpenStudy (anonymous):

Y=x-3, and y=-2x+3 , graph and find intersection point

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Y=x-3, and y=-2x+3 , graph and find intersection point

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

I doubt it

OpenStudy (anonymous):

Do the lines intersect ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i give up

OpenStudy (anonymous):

Do u have a graphing calculator

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

-.6, 1.8

OpenStudy (anonymous):

Is that a Option

OpenStudy (anonymous):

My mistake answer is -3 , 9

OpenStudy (anonymous):

Do yourself a favor and invest in a graphing calculator

OpenStudy (anonymous):

Ok u got it ??

OpenStudy (anonymous):

yes , thanks